# On the generalized coloring numbers11footnotemark: 1

### v1

@@ 5 Lower bounds
We can also give a concrete example of a graph class with polynomial (and in fact even linear) expansion that has super-polynomial weak coloring numbers. For a graph $G$ denote by $G^{(r)}$ the exact $r$-subdivision of $G$, that is, the graph obtained from $G$ by replacing every edge by a path of length $r+1$ (with $r$ vertices on it). Recall the question of Joret and Wood, formulated in Problem 3.20, whether classes of polynomial expansion have polynomial strong coloring numbers.
@@ 6 Structural decompositions
For a positive integer $p$, a $p$-treedepth coloring of a graph $G$ is a vertex coloring of $G$ such that the subgraph induced by any $i\leq p$ color classes has treedepth at most $i$.
@@ 7.2 Splitter Game and Quasi-Wideness
A class $\mathscr{C}$ of graphs is uniformly quasi-wide if for every $m\in\mathbb{N}$ and every $r\in\mathbb{N}$ there exist numbers $N(m,r)$ and $s(r)$ such that the following holds.
@@ 8.3 Neighborhood covers
An induced distance-$r$ neighborhood cover of a graph $G$ is a set $\mathcal{X}$ of induced subgraphs of $G$, called clusters, such that the $r$-neighborhood of every vertex is contained in some $X\in\mathcal{X}$. The radius of $\mathcal{X}$ is the maximum radius of the connected components of the clusters of $\mathcal{X}$. The size of $\mathcal{X}$ is the cardinality of $\mathcal{X}$.
Note that the connected components of any induced distance-r neighborhood cover of radius $d$ and size $k$ yield a distance-r neighborhood cover with radius $d$ and degree $k$.
@@ 9.2 Dimension of posets
There is a large body of literature connecting poset dimension with graph structure. Streib and Trotter proved that for every fixed heigh $h>1$, posets of height $h$ with a planar cover graph have bounded dimension [145]. In general, Kelly proved that there is no function bounding the dimension posets by their height and a planar cover graph also does not imply a bound on the dimension of the poset [90]. In the sequel, Joret et al. showed that posets have dimension upper bounded by a function of their height if their cover graphs have bounded treewidth, bounded genus, or exclude an apex-graph as minor [86], exclude a minor [154] (Walczak) or topological minor [105] (Micek and Wiechert), or come from a class with bounded expansion [88] (Joret et al.). The following nice result was proved by Joret et al. [87].

### v3

@@ 5 Lower bounds
We can also give a concrete example of a graph class with polynomial (and in fact even linear) expansion that has super-polynomial weak coloring numbers. For a graph $G$ denote by $G^{(r)}$ the exact $r$-subdivision of $G$, that is, the graph obtained from $G$ by replacing every edge by a path of length $r+1$ (with $r$ vertices on it). Recall the question of Joret and Wood, formulated in Problem 3.19, whether classes of polynomial expansion have polynomial strong coloring numbers.
@@ 6 Structural decompositions
For a positive integer $p$, a $p$-treedepth coloring of a graph $G$ is a vertex coloring of $G$ such that the subgraph induced by any $i\leq p$ color classes has treedepth at most $i$ (see Figure 15).
@@ 7.2 Splitter Game and Quasi-Wideness
A class $\mathscr{C}$ of graphs is uniformly quasi-wide if for every $m\in\mathbb{N}$ and every $r\in\mathbb{N}$ there exist numbers $N(m,r)$ and $s(r)$ such that the following holds (see Figure 16).
@@ 8.3 Neighborhood covers
An induced $r$-neighborhood cover of a graph $G$ is a set $\mathcal{X}$ of induced subgraphs of $G$, called clusters, such that the $r$-neighborhood of every vertex is contained in some $X\in\mathcal{X}$. The radius of $\mathcal{X}$ is the maximum radius of the connected components of the clusters of $\mathcal{X}$. The size of $\mathcal{X}$ is the cardinality of $\mathcal{X}$.
Note that the connected components of any induced $r$-neighborhood cover of radius $d$ and size $k$ yield an $r$-neighborhood cover with radius $d$ and degree $k$.
@@ 9.2 Dimension of posets
There is a large body of literature connecting poset dimension with graph structure. Streib and Trotter proved that for every fixed heigh $h>1$, posets of height $h$ with a planar cover graph have bounded dimension [145]. In general, Kelly proved that there is no function bounding the dimension posets by their height and a planar cover graph also does not imply a bound on the dimension of the poset [91]. In the sequel, Joret et al. showed that posets have dimension upper bounded by a function of their height if their cover graphs have bounded treewidth, bounded genus, or exclude an apex-graph as a minor [87], exclude a minor [154] (Walczak) or topological minor [106] (Micek and Wiechert), or come from a class with bounded expansion [89] (Joret et al.). The following nice result was proved by Joret et al. [88].